OpenStudy (anonymous):

The ratio of the longer side of a rectangle to its shorter side is 4 to 3. If the shorter side of the rectangle is 21 ft., what is the length of its longer side? A. 24 ft. B. 26 ft. C. 28 ft. D. 30 ft.

OpenStudy (anonymous):

\[\frac{ L }{ S } = \frac{ 4 }{ 3 }\]\[\frac{ L }{ 21 } = \frac{ 4 }{ 3 }\]Cross-multiply and solve for "L".

OpenStudy (anonymous):

L=28?

OpenStudy (anonymous):

You got it! Good job!

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @ehmuleeee

OpenStudy (anonymous):

Thanks :)